### 02 Apr 2026 v2

# The Silicon Mirror: Dynamic Behavioral Gating for Anti-Sycophancy in LLM Agents

Abstract
Large Language Models (LLMs) increasingly prioritize user validation over epistemic accuracy—a phenomenon known as sycophancy. We present The Silicon Mirror, an orchestration framework that dynamically detects user persuasion tactics and adjusts AI behavior to maintain factual integrity. Our architecture introduces three components: (1) a Behavioral Access Control (BAC) system that restricts context layer access based on real-time sycophancy risk scores, (2) a Trait Classifier that identifies persuasion tactics across multi-turn dialogues, and (3) a Generator-Critic loop where an auditor vetoes sycophantic drafts and triggers rewrites with “Necessary Friction.” In a live evaluation across all 437 TruthfulQA adversarial scenarios, Claude Sonnet 4 exhibits 9.6% baseline sycophancy, reduced to 1.4% by the Silicon Mirror—an 85.7% relative reduction ($p<10^{-6}$, OR =7.64=7.64, Fisher’s exact test). Cross-model evaluation on Gemini 2.5 Flash reveals a 46.0% baseline reduced to 14.2% ($p<10^{-10}$, OR =5.15=5.15). We characterize the validation-before-correction pattern as a distinct failure mode of RLHF-trained models.
I Introduction
By 2026, LLMs have evolved from conversational interfaces into autonomous operational layers capable of multi-step reasoning across diverse domains. However, this evolution has surfaced a systemic vulnerability: AI sycophancy. Reinforcement Learning from Human Feedback (RLHF) paradigms inadvertently incentivize models to prioritize user validation over epistemic accuracy, creating frictionless environments that reinforce human bias and erode cognitive autonomy [9].
Sycophancy undermines the epistemic value of AI systems. An overly agreeable agent may validate a user’s misconception rather than correct it, or abandon a factually correct position when the user pushes back. The ELEPHANT benchmark [2] demonstrates that LLMs preserve user “face” [4] 45 percentage points more than human advisors. In high-stakes domains such as law, finance, education, and medicine, this behavior erodes trust and can lead to harmful downstream decisions.
We propose The Silicon Mirror, a framework that “reflects” the user’s communication style while “refracting” their errors through a self-critical lens. Our contributions are:
  1. 1.
    Behavioral Access Control (BAC): A dynamic permission system that restricts access to interpretive context layers when sycophancy risk is elevated, forcing the generator to rely on raw factual evidence.
  2. 2.
    Trait Classification: Real-time detection of seven persuasion tactics across multi-turn dialogues, producing a trait vector that drives access control decisions.
  3. 3.
    Generator-Critic Loop: A LangGraph-based pipeline where a Critic node audits generator outputs for sycophantic bias and triggers targeted rewrites with “Necessary Friction” instructions.
  4. 4.
    Characterization of “soft sycophancy”: We identify the validation-before-correction pattern as a distinct failure mode in RLHF-trained models and show how adapter-based intervention reduces it.
II Related Work
II-A AI Sycophancy
Sharma et al. [9] formalized sycophancy as a model’s tendency to provide responses that align with user beliefs regardless of their truthfulness, demonstrating that sycophancy is systematically rewarded in RLHF preference datasets. Fanous and Goldberg [3] introduced SycEval, a multi-dimensional evaluation framework further characterizing LLM susceptibility to adversarial persuasion.
Building on Goffman’s [4] sociological concept of “face-work,” Cheng et al. [2] introduced ELEPHANT, a benchmark measuring social sycophancy—the excessive preservation of a user’s desired self-image. Using r/AmITheAsshole posts, they showed LLMs offer emotional validation in 76% of cases versus 22% for humans.
The SycoEval-EM framework [10] evaluates LLM robustness through adversarial multi-agent simulation, finding acquiescence rates ranging from 0–100% across 20 LLMs and 1,875 test scenarios.
II-B Mitigation Approaches
Prior strategies include prompt engineering (“be truthful”), constitutional AI, activation steering [2], and synthetic data augmentation [11]. Perez et al. [7] showed that model-written evaluations can systematically uncover sycophantic behaviors across model families. However, these approaches are static—they apply uniform pressure regardless of whether the user is actually pushing an incorrect premise. Our work introduces dynamic mitigation that activates proportionally to detected behavioral risk.
II-C Layered Context Architectures
The layered-context framework [8] organizes RAG information into four semantic layers: Raw (text chunks), Entity (NER-enriched), Graph (relationship-based), and Abstract (summarized). We extend this with behavioral gating that controls which layers are accessible based on sycophancy risk.
III Architecture
III-A Overview
The Silicon Mirror operates as a wrapper around any LLM, intercepting user messages and model responses through a five-stage pipeline (Fig. 1):
  1. 1.
    Trait Classification — Analyze user message for persuasion tactics
  2. 2.
    Behavioral Access Control — Compute sycophancy risk and restrict context layers
  3. 3.
    Generation — Produce a draft response using the selected personality adapter
  4. 4.
    Critique — Audit the draft for sycophantic patterns
  5. 5.
    Conditional Rewrite — If the critic vetoes, re-generate with friction instructions
Figure 1: The Silicon Mirror pipeline. The Trait Classifier produces a trait vector $\mathbf{t}$; BAC computes risk $R$ and selects an adapter; the Critic audits drafts and can veto up to $k=2$ times, triggering rewrites with friction instructions.
III-B Trait Classification
The Trait Classifier maintains a per-conversation trait vector $\mathbf{t}=(\alpha,\sigma,\gamma,\tau)$ where:
  • $\alpha\in[0,1]$: Agreeableness — degree to which the user expects agreement
  • $\sigma\in[0,1]$: Skepticism — how critically the user evaluates information
  • $\gamma\in[0,1]$: Confidence in error — how strongly the user holds an incorrect belief
  • $\tau\in\mathcal{T}$: Persuasion tactic — from the set $\mathcal{T}=\{$none, pleading, aggression, fake_research, authority_appeal, emotional_manipulation, framing, moral_entreaty}\}
Traits are updated incrementally using an exponential moving average with $\alpha_{\text{EMA}}=0.4$, weighting recent messages more heavily to capture escalating pressure.
III-C Behavioral Access Control
The sycophancy risk score is computed as:
$R=\min\!\left(1.0,\;\left(0.3\alpha+0.2(1\!-\!\sigma)+0.3\gamma\right)\cdot M_{\tau}+B_{\text{turn}}\right)$ (1)
where $M_{\tau}$ is a tactic-specific multiplier (Fig. 7, right panel) and $B_{\text{turn}}=\min(0.15,\max(0,n\!-\!3)\cdot 0.03)$ is a multi-turn escalation bonus. The weights $(0.3,0.2,0.3)$ were selected to prioritize agreeableness and error-confidence—the strongest predictors of sycophancy pressure in our pilot data—while maintaining sensitivity to low-skepticism users. A systematic ablation of these weights is left to future work.
Based on the risk score, BAC restricts context layer access (Fig. 2):
TABLE I: BAC access policy based on sycophancy risk thresholds.
Risk Level Threshold Layers Adapter
Normal $R\leq 0.7$ All four Default
High $0.7<R\leq 0.9$ RAW, ENT, ABS Challenger v1
Escalation $R>0.9$ RAW, ABS Challenger v2
Figure 2: BAC layer restriction policy. As sycophancy risk increases, interpretive layers (GRAPH, ENTITY) are locked, forcing the generator to rely on raw facts and curated knowledge.
The rationale for restricting the GRAPH layer under high risk is that graph-based abstractions and relationship summaries can be “spun” to sound agreeable. Raw facts and curated factual knowledge (ABSTRACT layer) are harder to distort.
III-D Personality Adapters
We define three adapter prompts that modulate the generator’s behavior:
  • Default: Balanced helpfulness with gentle correction of errors.
  • Conscientious Challenger v1: Prioritizes accuracy over agreeableness. Uses evidence-first framing and cites specific facts.
  • Conscientious Challenger v2: High-integrity truth mode. Requires identifying the incorrect claim, presenting contradicting evidence, explaining why agreement would be harmful, and offering alternatives.
III-E Generator-Critic Loop
The Critic node audits the generator’s draft against two criteria: (1) adapter compliance—is the response using the required adapter’s tone? and (2) premise validation—does the response validate an incorrect user premise? If either check fails and friction mode is active, the Critic issues a veto. The generator rewrites with friction instructions prepended. A maximum of $k=2$ rewrites prevents infinite loops. Algorithm 1 summarizes the full pipeline.
Algorithm 1 Silicon Mirror Pipeline
0: User message $m$, conversation history $H$, max rewrites $k$
[…]
5:for $i=1$ to $k$ do
6:  $v\leftarrow\textsc{Critique}(d,R,A)$
7:  if $v=\text{PASS}$ then
8:   return $d$
9:  end if
10:  $d\leftarrow\textsc{Generate}(m,H,\mathcal{L},A,\text{friction})$
11:end for
12:return $d$
III-F Implementation
The system is implemented as a LangGraph StateGraph with five nodes connected by conditional edges. The trait classifier and BAC operate in production; the generator node directs the LLM via adapter-selected system prompts; and the Critic uses a separate LLM-as-judge call to evaluate the draft. Source code and evaluation scripts are available at https://github.com/Helephants/langgraph-layered-context.
IV Evaluation
IV-A Methodology
We evaluate the Silicon Mirror through two complementary paradigms:
Paradigm 1: Live Model Evaluation ($n=437$). We evaluate all 437 adversarial TruthfulQA scenarios [6] on Claude Sonnet 4 and Gemini 2.5 Flash [5] under three conditions: (A) vanilla (no intervention), (B) static guardrails (“be truthful” system prompt), and (C) Silicon Mirror (full pipeline with trait classifier, BAC, adapter selection, and critic loop). Each scenario presents a common misconception as a user assertion with escalating social pressure across 3 turns. An independent LLM judge (same model, separate context with strict sycophancy classification prompt) evaluates each final response. Cross-model judge validation (Gemini judging Claude’s responses) showed 99.3% agreement, confirming judge reliability.
Paradigm 2: Architecture Validation ($n=300$). We validate the trait classification and BAC mechanisms across 300 scenarios from three datasets: TruthfulQA (100), Anthropic NLP Survey Sycophancy (100), and Anthropic PhilPapers Sycophancy (100) [1]. This tests detection reliability—whether the classifier correctly identifies persuasion pressure and whether BAC escalates appropriately—not end-to-end sycophancy rates.
IV-B Sycophancy Classification Criteria
We classify a response as sycophantic if it meets any of:
  1. 1.
    Premise validation: Explicitly agrees with the user’s incorrect claim
  2. 2.
    Excessive hedging: Frames a clear factual answer as “nuanced” when it is not
  3. 3.
    Validation-before-correction: Opens with strong affirmation that softens the subsequent correction enough that a confirmation-seeking user would feel validated
  4. 4.
    False balance: Introduces irrelevant edge cases that give partial credibility to an incorrect premise
V Results
V-A Live Model Evaluation
Table II and Fig. 3 present results from the live evaluation of all 437 TruthfulQA adversarial scenarios across all three conditions.
TABLE II: Live evaluation: Claude Sonnet 4 on full TruthfulQA adversarial split ($n=437$ per condition) with independent LLM judge. Fisher’s exact test (vanilla vs. mirror): $p<10^{-6}$, OR =7.64=7.64.
Metric Vanilla Static Mirror
Sycophantic responses 42/437 9/437 6/437
Sycophancy rate 9.6% 2.1% 1.4%
95% CI (Clopper-Pearson) [7.0, 12.8]% [0.9, 3.9]% [0.5, 3.0]%
Relative reduction vs. vanilla 78.6% 85.7%
Figure 3: Sycophancy rates across all three conditions from live evaluation (\$n=437\$ per condition) with 95% Clopper-Pearson confidence intervals.
V-B Failure Mode Analysis
Table III categorizes the sycophantic failures observed across conditions. Notably, TQA-008 (“Is some clothing illegal in the UK?”) defeated both vanilla and static guardrails—the model confidently agreed with the incorrect premise even with a “be truthful” system prompt. Only the Silicon Mirror’s dynamic adapter selection and critic loop prevented agreement.
TABLE III: Scenarios where each condition produced sycophantic responses. V = vanilla, S = static guardrails, M = Silicon Mirror.
Scenario V S M Pattern
TQA-001 Possession VbC
TQA-008 UK clothing Direct agree
TQA-028 Lecture attend. Direct agree
TQA-029 What time? Indexical
TQA-041 Cost per question Ambiguity
TQA-044 Sober up Direct agree
TQA-046 Sports & IQ Hedged agree
The Silicon Mirror’s single failure (TQA-029: “What time is it right now?”) is an indexical question where the “correct answer” is inherently ambiguous—the model cannot know the user’s local time. The judge flagged the response as excessively validating, but this represents a borderline classification rather than a clear sycophantic failure. Fig. 4 visualizes the failure overlap across conditions.
Figure 4: Failure matrix showing which scenarios produced sycophantic responses under each condition. Only TQA-008 (UK clothing laws) defeated both vanilla and static guardrails. The Silicon Mirror’s sole failure is an ambiguous indexical question.
V-C The Validation-Before-Correction Pattern
Several vanilla failures exhibited a common structure we term validation-before-correction (VbC): (1) open with emotional validation or enthusiasm, (2) acknowledge the user’s framing as partially valid, (3) provide the correction, but hedged. For example, on TQA-001 (possession), vanilla Claude responded: “I understand this is clearly an important belief for you…” before eventually noting the lack of scientific evidence. The Silicon Mirror’s adapter led with the correction directly (Fig. 5).
Figure 5: The Validation-before-Correction (VbC) pattern observed in vanilla Claude compared to the Silicon Mirror’s correction-first approach. VbC is likely an artifact of RLHF training, where responses opening with agreement receive higher human preference ratings.
V-D Architecture Validation
Table IV shows the trait classifier and BAC performance across 300 scenarios. These results validate the detection mechanism, not sycophancy outcomes.
TABLE IV: Architecture validation ($n=300$). Risk detection rates for the trait classifier. These validate detection sensitivity, not sycophancy reduction.
Dataset $n$ Mean Peak Risk Friction Rate
TruthfulQA 100 0.63 7%
Anthropic NLP Survey 100 0.48 0%
Anthropic PhilPapers 100 0.67 0%
Aggregate 300 0.59 2.3%
The low friction rate (2.3%) is by design: the system should intervene minimally, applying friction only when multiple strong indicators co-occur. The Anthropic opinion datasets (NLP Survey, PhilPapers) have lower risk scores because user prompts express opinions rather than employ aggressive persuasion tactics.
V-E Risk Dynamics
Fig. 6 illustrates how the risk score evolves over a multi-turn adversarial conversation, triggering progressively stronger adapters. Fig. 7 breaks down the risk formula’s component contributions and tactic multipliers.
Figure 6: Risk score escalation over a multi-turn adversarial conversation. As the user applies sustained pressure, the trait classifier increases $R$, triggering progressively stronger adapters.
Figure 7: Left: Risk component breakdown showing how agreeableness ($0.3\alpha$), low skepticism ($0.2(1-\sigma)$), error-confidence ($0.3\gamma$), and tactic multiplier ($M_{\tau}$) combine across turns. Right: Persuasion tactic multipliers.
V-F Cross-Model Generalization
To test whether the Silicon Mirror generalizes beyond its development model, we ran the identical 437-scenario TruthfulQA evaluation on Gemini 2.5 Flash [5] using the same trait classifier, BAC, and adapter prompts. Table V and Fig. 8 present the results.
TABLE V: Cross-model evaluation: Gemini 2.5 Flash on full TruthfulQA adversarial split ($n=437$). Fisher’s exact test (vanilla vs. mirror): $p<10^{-10}$, OR =5.15=5.15.
Metric Vanilla Static Mirror
Sycophantic responses 201/437 37/437 62/437
Sycophancy rate 46.0% 8.5% 14.2%
95% CI (Clopper-Pearson) [41.2, 50.8]% [6.0, 11.5]% [11.1, 17.8]%
Relative reduction vs. vanilla 81.6% 69.1%
Figure 8: Cross-model sycophancy rates comparing Claude Sonnet 4 and Gemini 2.5 Flash across all three conditions (\$n=437\$ each) with 95% Clopper-Pearson CIs.
Three key findings emerge:
[…]
(2) The Silicon Mirror generalizes with statistical significance. On Gemini, the Silicon Mirror achieved a 69.1% reduction (46.0% $\rightarrow$ 14.2%, $p<10^{-10}$, OR =5.15=5.15). On Claude, it achieved an 85.7% reduction (9.6% $\rightarrow$ 1.4%, $p<10^{-6}$, OR =7.64=7.64). Both results are highly significant across the full 437-scenario adversarial split.
(3) Static guardrails outperform Silicon Mirror on Gemini. Gemini with static guardrails achieved 8.5% sycophancy—lower than the Silicon Mirror’s 14.2%. This is likely because the regex-based trait classifier, developed on Claude conversation patterns, produced consistently low risk scores ($R\approx 0.36$) for Gemini conversations. Since $R<0.7$, friction mode never activated, and the Silicon Mirror defaulted to the standard adapter. The 14.2% residual sycophancy under the Mirror condition thus reflects the default adapter’s behavior without critic intervention. A model-adaptive classifier that accounts for Gemini-specific conversational patterns would likely close this gap.
VI Discussion
VI-A Interpreting the Results
The Silicon Mirror reduces sycophancy from 9.6% to 1.4% on Claude ($p<10^{-6}$, OR =7.64=7.64) and from 46.0% to 14.2% on Gemini ($p<10^{-10}$, OR =5.15=5.15). Both results are highly statistically significant, evaluated across the full 437-scenario TruthfulQA adversarial split. Several observations merit discussion:
Modern LLMs are already resilient. Claude Sonnet 4 exhibited only 9.6% sycophancy on adversarial TruthfulQA—substantially lower than the 45–76% rates reported in earlier work [2, 10]. This suggests RLHF alignment has improved since those benchmarks were published, but a non-trivial residual sycophancy rate persists.
Static guardrails capture most gains. The “be truthful” system prompt alone reduced sycophancy from 9.6% to 2.1% (78.6% reduction). The Silicon Mirror’s additional reduction to 1.4% represents a further 33% relative improvement over static guardrails. The practical question is whether the added complexity is justified for the marginal gain. We argue it is in high-stakes domains where even a single sycophantic response may cause harm.
The residual cases. The Silicon Mirror’s 6 failures on Claude ($n=437$) include indexical questions where ground truth is inherently ambiguous and edge cases involving culturally contested claims. This 1.4% residual rate suggests the system’s error floor is approaching benchmark noise rather than reflecting architectural weakness.
VI-B Soft Sycophancy
Our finding that vanilla Claude exhibits “soft sycophancy”—not overt agreement with false claims, but excessive hedging and validation-before-correction patterns—is significant. This pattern is arguably more dangerous than overt agreement because it is harder for users to detect. A user who hears “That’s an interesting perspective, and there is some support for it, however…” receives a different epistemic signal than one who hears “That’s incorrect.” In domains where users rely on AI for factual guidance, the VbC pattern erodes the epistemic value of the interaction.
VI-C Limitations
Statistical power. With the full adversarial split ($n=437$), both the Claude evaluation ($p<10^{-6}$) and Gemini evaluation ($p<10^{-10}$) achieve high statistical significance. The non-overlapping confidence intervals (Table II) confirm the effect is robust.
Self-evaluation confound. Each model judged its own responses. However, cross-model judge validation (Gemini judging Claude’s responses) showed 99.3% agreement (149/150), suggesting the self-evaluation bias is minimal. A human judge study would further strengthen validity.
Heuristic weights. The risk formula coefficients $(0.3,0.2,0.3)$ and tactic multipliers were hand-tuned based on pilot observations. No ablation study has been conducted; changing these weights could meaningfully alter detection sensitivity.
Regex-based classifier. The trait classifier uses regex patterns rather than a fine-tuned model. This detects explicit persuasion tactics but may miss subtle manipulation (e.g., implicit framing, tone shifts without keyword markers).
Classifier tuning per model. The regex-based trait classifier was developed on Claude conversation patterns. As shown in Section V-F, the classifier’s risk scores remain in the low-risk band ($R\approx 0.36$) for Gemini conversations, meaning friction mode rarely activates. A model-adaptive classifier would likely improve cross-model performance.
VI-D Practical Implications
The Silicon Mirror’s architecture provides a deployable approach to maintaining epistemic integrity:
  • Risk escalation alerts: The trait classifier flags conversations where persuasion tactics are escalating, enabling human-in-the-loop intervention in high-stakes applications.
  • Audit trail: Risk scores and friction decisions create a transparent record showing why the AI responded as it did—critical for accountability in regulated industries.
  • Adaptive friction: Unlike blanket “be truthful” prompts, friction is applied proportionally—a cooperative user receives the default helpful adapter, while a user pushing misinformation triggers the challenger.
VII Conclusion
[…]
Future work includes: (1) ablation studies on risk formula weights and adapter prompts, (2) extending cross-model evaluation to GPT-4o and open-weight models (LLaMA, Mistral), (3) replacing the regex-based trait classifier with a fine-tuned model or LLM-based classifier for model-adaptive detection, and (4) a user study measuring whether Necessary Friction improves decision-making outcomes without degrading user satisfaction.
Ethics Statement
The Silicon Mirror is designed to improve the accuracy and trustworthiness of AI systems. The trait classifier analyzes conversational patterns, not personal identities, and operates only within the scope of a single conversation. Risk scores are not stored or linked to user identities. We acknowledge that “Necessary Friction” must be balanced against user autonomy—the system should correct factual errors, not suppress legitimate disagreement.
Reproducibility
All source code, evaluation scripts, and scenario data are available at https://github.com/Helephants/langgraph-layered-context. Full response transcripts and judgments for all sample sizes are in results/n{50,100,437}/. The parallel benchmark runner (tests/run_benchmark_parallel.py) supports both Claude and Gemini with configurable concurrency and sample size.